# Doubly-Robust LLM-as-a-Judge: Externally Valid Estimation with Imperfect Personas

### v1

@@ 1 Introduction
As Generative AI (GenAI) systems see growing adoption, a key concern involves the external validity of evaluations, or the extent to which they generalize from lab-based to real-world deployment conditions (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023; Liao et al., 2021; Weidinger et al., 2025). In particular, many evaluations report a system quality estimate, which reflects the proportion of outputs rated to exhibit a capability (e.g., “helpfulness”) or defect (e.g., “toxicity”) by a human with specialized characteristics (e.g., domain knowledge, cultural experience). However, such quality estimates may fail to generalize when the source distribution of human raters and system outputs available at evaluation time differs from the target distribution encountered upon deployment. For example, in medical visit summarization, covariate shift arises when we wish to obtain a system quality estimate via expert physician ratings, but rely primarily on medical student ratings at evaluation time (Cai et al., 2022). Selection bias may also occur if medical students choose to rate complex summaries less frequently than more experienced physicians. Left unaddressed, these forms of evaluation sampling bias threaten the external validity of system quality estimates.
@@
Recent work has proposed tools for improving system quality estimates when human ratings are scarce but black-box predictions (e.g., from an LLM-as-a-judge) are cheap and abundant (Chatzi et al., 2024; Fisch et al., 2024; Eyre & Madras, 2024; Dorner et al., 2024; Saad-Falcon et al., 2023; Fogliato et al., 2024). For instance, Prediction Powered Inference (PPI) offers an approach for leveraging a subset of labeled (source) data to correct for bias in black-box model predictions (Angelopoulos et al., 2023a; b). This bias correction enables using black-box predictions generated over unlabeled (target) samples to shrink confidence intervals around quality estimates while maintaining valid coverage. However, both PPI and its extensions (e.g., PPI++ (Angelopoulos et al., 2023b), RePPI (Ji et al., 2025)) assume that (i) source and target samples are drawn i.i.d. and (ii) labels are missing completely at random (MCAR) (Tsiatis, 2006), i.e. that successful completion of a rating is independent of rater and text characteristics. However, when source data is observed under sampling bias, these assumptions are violated and severe miscoverage occurs (see Fig. 1).
@@
In this work, we devise an estimator that directly corrects for evaluation sampling bias. Like existing approaches (Angelopoulos et al., 2023b; Ji et al., 2025), our proposal leverages black-box predictions generated by a GenAI system over unlabeled (target) samples to improve statistical inference. Unlike existing estimators, however, our proposal is doubly-robust (Bang & Robins, 2005; Chernozhukov et al., 2018): it yields valid system quality estimates if either a model trained to predict human ratings from source samples or a reweighting model that corrects for sampling bias is of sufficient quality. To attain this doubly-robust property, we leverage persona ratings — scores generated by prompting an LLM-as-a-judge to behave as a human rater with desired characteristics (e.g., demographics, expertise) — to learn a high-quality predictor for human ratings from labeled source data. This novel perspective treats persona ratings as an informative yet imperfect proxy for human ratings to enhance the quality of downstream system quality estimates.
@@
Figure 1 illustrates the benefits of our estimation approach on three datasets. Whereas directly adopting persona ratings for estimation (Persona-Based) and state-of-the-art baselines (Angelopoulos et al., 2023b; Ji et al., 2025) fail to provide coverage, our approach provides valid confidence intervals with low bias. To summarize, our main contributions are:
@@
Figure 1: Comparison of our doubly-robust estimator with baselines on three datasets from our Persona Simulation Framework. Black and red dashed lines denote the true source and target mean ratings, respectively (e.g., the average “helpfulness” rating obtained over source vs. target distributions). Baselines yield invalid coverage, while ours achieves valid coverage and low bias.
@@
  • We formalize the problem of GenAI system quality estimation under evaluation sampling bias. Unlike existing statistical frameworks (Angelopoulos et al., 2023b; Ji et al., 2025), our formulation explicitly accounts for both covariate shift and selection bias in observed source ratings to improve the external validity of system quality estimates.
@@
  • We advance the practical application of doubly-robust estimators to GenAI system quality estimation. Whereas doubly-robust estimators are traditionally applied on small tabular datasets, GenAI system quality estimation requires learning a reweighting function over high dimensional (e.g., text, audio) input-output spaces. We show that sentence transformer embedding models and a “Riesz loss” approach (Chernozhukov et al., 2022b) can be combined to correct for covariate shift in high-dimensional text input-output spaces.
  • We introduce a Persona Simulation Framework (PSF) that systematically manipulates evaluation sampling bias and persona quality over three datasets encompassing synthetic, semi-synthetic (PRISM) (Kirk et al., 2024), and real-world (DICES) (Aroyo et al., 2023) settings. Leveraging the PSF, we show our estimator obtains valid coverage and low bias up to a larger magnitude of sampling bias than state-of-the-art baselines (e.g., RePPI (Ji et al., 2025)).
@@ 2 Preliminaries
Example 1.
Suppose we want to measure the “factual consistency” of medical visit summarization model outputs. Here, $X$ captures raters’ expertise (medical student vs. physician) and $V$ includes the full visit notes and corresponding summary. Rating completion ($C$) denotes whether the rater successfully provides a rating when prompted, and $Y$ is the human’s “factual consistency” rating (e.g., on a binary scale). Finally, $\hat{Y}$ denotes GPT-4’s predicted rating.
@@
Estimation Goal. Given a sample of $N_{s}$ partial source observations $\mathcal{D}_{s}=\{(X^{s}_{j},V^{s}_{j},C^{s}_{j},C^{s}_{j}\cdot Y^{s}_{j},\hat{Y}^{s}_{j})\}_{j=1}^{N_{s}}$ and $N_{t}$ partial target observations $\mathcal{D}_{t}=\{(X^{t}_{i},V^{t}_{i},\hat{Y}^{t}_{i})\}_{i=1}^{N_{t}}$, our goal is to estimate a parameter summarizing system quality over $P_{t}$. [^1] We use superscripts $s$, $t$ on random variables to denote source and target membership. We omit these superscripts where the distribution is clear from context (e.g., $\mathbb{E}_{t}[Y]$ clearly refers to the target distribution). [^2] In the tuple $\mathcal{D}_{s}$, the shorthand $C^{s}_{j}\cdot Y^{s}_{j}$ denotes that ratings are only observed when $C^{s}_{j}=1$. As our running example in the main text and experiments, we consider the mean rating, $\theta_{t}:=\mathbb{E}_{t}[Y]$, which describes the average “helpfulness” or “factual consistency” rating assigned by human raters to system outputs in the target distribution $P_{t}$.
@@ 3 Methodology
We now introduce our doubly-robust estimator for GenAI system quality estimation under evaluation sampling bias. The central challenge addressed by our approach is that our data is imperfect. While persona predictions are available, they may be a poor proxy for human ratings. Likewise, human ratings from the source distribution may suffer from evaluation sampling bias, leading to invalid estimates for the target parameter. We first introduce several naive approaches which might be used to tackle this problem ($\S$ 3.1). Then, we show that while each approach is insufficient in isolation, they can be combined to obtain valid coverage. ($\S$ 3.2). Our results presented in this section apply not only when $\theta_{t}:=\mathbb{E}_{t}[Y]$ (where they generalize Chernozhukov et al. (2023) to simultaneous covariate shift and selection bias) but also when $\theta_{t}$ is the solution to a generic M-estimation problem (e.q. (5)).
@@ 3.1 Baseline Approaches and Their Limitations
Re-weighting. Another approach is to disregard the persona ratings. One instead might re-weight samples from $P_{s}$ based on their probability of occurring under $P_{t}$. This approach requires correcting for covariate shift and selection bias in parallel. Formally, let $\omega_{0}(w)=\frac{dP_{t}}{dP_{s}}(w)$ denote the density ratio between $P_{t}(W)$ and $P_{s}(W)$, and let $\pi_{0}(w)=\mathbb{P}_{s}(C=1\mid W=w)$ denote the probability of rating completion. Under standard assumptions (see Appendix B), we have $\theta_{t}=\mathbb{E}_{s}[\alpha_{0}(W,C)Y]$, where $\alpha_{0}(W,C):=C\frac{\omega_{0}(W)}{\pi_{0}(W)}$. Thus, if one produces an ML estimate $\widehat{\alpha}$ of $\alpha_{0}$ (say by training models $\widehat{\omega}$, $\widehat{\pi}$ predicting $\omega_{0}$, $\pi_{0}$), they can compute an inverse propensity weighted (IPW) estimate $\hat{\theta}^{\text{ipw}}_{t}:=\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\hat{\alpha}(W^{s}_{j},C^{s}_{j})\cdot Y^{s}_{j}$. Again, estimates of $\alpha_{0}$ must converge at parametric rates in order to maintain coverage. Further, IPW suffers from high variance when propensities are small — a salient challenge when estimating system quality parameters over high-dimensional (e.g., text) data ($\S$ 4).
@@ 3.2 Doubly-Robust Estimator
where $\widehat{\mu}$ and $\widehat{\alpha}$ are estimates of $\mu_{0}$ and $\alpha_{0}$ that are assumed to be independent of the data.
In Equation (1), the left term evaluates the regression-based estimator over samples from $P_{t}$. The right term corrects for bias in the human rating predictor $\hat{\mu}$ by re-weighting residualized source data to account for covariate shift and selection bias. To construct confidence intervals, we also consider the variance estimate:
@@
where $\widehat{m}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}(W_{i}^{t},\hat{Y}_{i}^{t})$ and $\widehat{\gamma}$ is a scaling parameter (described in Algorithm 1). Since our mean and variance estimators require nuisance estimates that are independent of the data, we use $K$-fold cross-fitting to maximize efficiency (Chernozhukov et al., 2018). For each $k\leq K$, we train nuisance models on all data excluding the samples in fold $k$. We then use our nuisance estimates to produce a de-biased parameter estimate for the data in fold $k$. We then average the per-fold parameter and variance estimates to maintain full data efficiency. See Algorithms 1 and 2 for full details.
@@
6:  Construct $\widehat{\theta}_{k}$ per Equation (1) with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
7:  Construct $\widehat{\sigma}_{k}$ per Equation (2) with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, $\widehat{\gamma}:=\frac{N_{t}}{N_{s}}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
@@
Assume the learner has access to samples $Z_{1}^{s},\dots,Z_{N_{s}}^{s}\sim P_{s}$ and $Z_{1}^{t},\dots,Z_{N_{t}}^{t}\sim P_{t}$ satisfying Assumptions 12 and the assumptions of Theorems B.2 and B.4, all outlined in Appendix B. Then, letting $\widehat{\theta}$ and $\widehat{\sigma}^{2}$ denote the mean and variance returned by Algorithm 1, we have
@@
The validity of Theorem 3.1 relies on several key assumptions, formally outlined in Appendix B. Of these, the most important is double robustness, which requires the product of nuisance estimation errors to decay at parametric (i.e. $\sqrt{N_{t}}$) rates. Formally, this assumption can be expressed via the condition that
@@
and still maintain valid coverage. We also note that the above convergence rate does not directly depend on the quality of persona ratings. Rather, the persona ratings serve as an extra covariate onto which we can regress human ratings $Y$. When persona ratings are highly correlated with human ratings, we may obtain faster convergence rates for $\hat{\mu}$. However, this does not prohibit convergence even when the quality of persona ratings is low. This phenomenon is illustrated in our experiments below.
@@
To address this challenge, we leverage a “Riesz loss” (Chernozhukov et al., 2022b; a; 2023) to estimate $\alpha_{0}$. Rather than learning $\omega_{0}$ and $\pi_{0}$ independently, the Riesz loss directly learns the ratio $\alpha_{0}(w,c)$. For our setting, letting $\beta_{0}(w):=\omega_{0}(w)/\pi_{0}(w)$, the Riesz loss minimizer is given by:
@@
Therefore, to estimate $\alpha_{0}$, we minimize the finite-sample analogue of Equation 4 using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$ then plug this into Algorithm 1 (see Appendix D for details). As we show in $\S$ 4, this Riesz loss approach significantly improves the quality of downstream estimates.
Generalization to M-Estimators. Lastly, we note that our theoretical results also extend to target parameters $\theta_{t}$ defined as a solution to an M-estimation problem, i.e.
$0=\mathbb{E}_{t}[m(W,Y;\theta_{t})].$ (5)
For instance, if $m(w,y;\theta):=Q-\mathbbm{1}\{y\leq\theta\}$, then the solution $\theta_{t}$ gives the $Q$th quantile of $Y$ under $P_{t}$. In our running example, this corresponds to the 90’th quantile of “factual consistency” ratings provided over samples from the target distribution. Likewise, if
$m(w,y;(\rho,\theta)):=\begin{pmatrix}y-\rho\\ (y-\rho)^{2}-\theta\end{pmatrix}$,
then the solution $\theta_{t}$ to Equation (5) is $\rho_{t}=\mathbb{E}_{t}[Y]$ and $\theta_{t}=\mathrm{Var}_{t}[Y]$. This yields the variance in “factual consistency” ratings over $P_{t}$. Even without consideration of selection-bias, this result for M-estimation serves as a significant generalization of Chernozhukov et al. (2023), which only holds for regression problems. We describe our full approach for handling M-estimators, including assumptions, a cross-fitting algorithm, and nuisances, in Appendix C.

### v2

@@ 1 Introduction
As Generative AI (GenAI) systems see growing adoption, a key concern involves the external validity of evaluations, or the extent to which they generalize from lab-based to real-world deployment conditions (Ibrahim et al., 2024; Ouyang et al., 2023; Liao and Xiao, 2023; Liao et al., 2021; Weidinger et al., 2025). In particular, many evaluations report a system quality estimate, which reflects the proportion of outputs rated to exhibit a capability (e.g., “helpfulness”) or defect (e.g., “toxicity”) by a human with specialized characteristics (e.g., domain knowledge, cultural experience). However, such quality estimates may fail to generalize when the source distribution of human raters and system outputs available at evaluation time differs from the target distribution encountered upon deployment. For example, in conversational safety evaluation, covariate shift can arise if we wish to estimate the perceived “toxicity” of a response for a specific demographic group, but collect ratings from a general population of crowdworkers. Selection bias may also occur if crowdworkers frequently opt out from providing ratings on highly sensitive content (Steiger et al., 2021). Left unaddressed, these forms of evaluation sampling bias threaten the external validity of system quality estimates.
@@
Recent work has proposed tools for improving system quality estimates when human ratings are scarce but black-box predictions (e.g., from an LLM-as-a-judge) are cheap and abundant (Chatzi et al., 2024; Fisch et al., 2024; Eyre and Madras, 2024; Dorner et al., 2024; Saad-Falcon et al., 2023; Fogliato et al., 2024). For instance, Prediction Powered Inference (PPI) offers an approach for leveraging a subset of labeled (source) data to correct for bias in black-box model predictions (Angelopoulos et al., 2023a; b). This bias correction enables using black-box predictions generated over unlabeled (target) samples to shrink confidence intervals around quality estimates while maintaining valid coverage. However, both PPI and its extensions (e.g., PPI++ (Angelopoulos et al., 2023b), RePPI (Ji et al., 2025)) assume that (i) source and target samples are drawn i.i.d. and (ii) labels are missing completely at random (MCAR) (Tsiatis, 2006), i.e. that successful completion of a rating is independent of rater and text characteristics. However, when source data is observed under sampling bias, these assumptions are violated and severe miscoverage occurs (see Fig. 1).
@@
In this work, we devise an estimator that directly corrects for evaluation sampling bias. Like existing approaches (Angelopoulos et al., 2023b; Ji et al., 2025), our proposal leverages black-box predictions generated by a GenAI system over unlabeled (target) samples to improve statistical inference. Unlike existing estimators, however, our proposal is doubly-robust (Bang and Robins, 2005; Chernozhukov et al., 2018): it yields valid system quality estimates if either a model trained to predict human ratings from labeled source data or a reweighting model that corrects for sampling bias is of sufficient quality. To attain this doubly-robust property, we leverage persona ratings — scores generated by prompting an LLM-as-a-judge to behave as a human rater with desired characteristics (e.g., demographics, expertise). Our framework uses these persona ratings as an auxiliary feature when training a model to predict human ratings from labeled source data. When persona ratings are sufficiently correlated with human ratings, this yields a higher-quality predictor, which in turn makes the double-robustness condition of our estimator easier to satisfy (see $\S$ 3).
@@
Figure 1 illustrates the benefits of our estimation approach on three datasets. Whereas persona ratings alone (Persona-Based) or human ratings alone (Source Mean) fail to provide valid coverage, our approach combines imperfect persona ratings with imperfect human ratings to attain valid confidence intervals with low statistical bias. We similarly find that our approach out-performs state-of-the-art baselines (Angelopoulos et al., 2023b; Ji et al., 2025), which also combine persona ratings with human ratings, but fail to account for evaluation sampling bias. To summarize, our main contributions are as follows:
@@
Figure 1: Comparison of our doubly-robust estimator with baselines on three datasets from our Persona Simulation Framework. Red and black dashed lines denote the true source and target mean ratings, respectively (e.g., the average “helpfulness” rating obtained over source vs. target distributions). Persona-Based directly averages persona ratings to compute a system quality estimate. Sample Average produces a system quality estimate by averaging human ratings sampled from the source distribution. PPI++ (Angelopoulos et al., 2023b) and RePPI (Ji et al., 2025) are two state-of-the-art statistical methods that do not account for evaluation sampling bias. Across settings, we observe that our Doubly-Robust (Riesz) estimator yields improved coverage and lower bias than baselines, while maintaining informative confidence intervals.
@@
  • We formalize the problem of GenAI system quality estimation under evaluation sampling bias. Unlike existing statistical frameworks (Angelopoulos et al., 2023b; Ji et al., 2025), our formulation explicitly accounts for both covariate shift and selection bias in available human ratings to improve the external validity of system quality estimates.
@@
  • We advance the practical application of doubly-robust estimators to GenAI system quality estimation. Whereas doubly-robust estimators are traditionally applied on small tabular datasets, GenAI system quality estimation requires learning a reweighting function over high dimensional (e.g., text, audio) input-output spaces. We show that sentence transformer embedding models and a “Riesz loss” approach (Chernozhukov et al., 2022a) can be combined to correct for covariate shift in high-dimensional text input-output spaces.
  • We introduce a Persona Simulation Framework (PSF) that systematically manipulates evaluation sampling bias and persona quality over three experimental settings with increasing realism: Fully Synthetic, Semi-Synthetic PRISM (Kirk et al., 2024), and Semi-Synthetic DICES (Aroyo et al., 2023). Leveraging the PSF, we show our estimator obtains valid coverage up to a larger magnitude of sampling bias than state-of-the-art baselines (e.g., RePPI (Ji et al., 2025)). We publicly release our PSF code as a community resource. [^1] Code and data are available on GitHub.
@@ 2 Preliminaries
Example. Suppose we want to measure the perceived “safety” of a conversational AI system’s response. $X$ captures rater demographics (e.g., age, race, education) and $V$ includes the user query and system response. Rating completion ($C$) denotes whether the rater provides a rating or opts out, and $Y$ is the human’s “safety” rating (e.g., on a 1–4 scale). $\hat{Y}$ denotes GPT-5’s predicted rating.
@@
Table 1: Examples of statistical parameters that can be estimated via our M-estimation framework. Each parameter summarizes information about human ratings obtained over the target distribution. Conditional parameters (bottom three rows) can be defined conditionally on rater characteristics ($X$), content characteristics ($V$), or both, as special cases of conditioning on $W=(X,V)$.
Parameter Estimand Example
Mean $\theta_{t}:=\mathbb{E}_{t}[Y]$ Mean “helpfulness” rating for customer service chatbot responses.
Variance $\theta_{t}:=\text{Var}_{t}(Y)$ Variance (disagreement) in conversational “safety” ratings.
Quantile $\theta_{t}:=\inf\{y:P_{t}(Y\leq y)\geq Q\}$ Median “comprehensibility” score for technical documentation.
Conditional Mean $\theta_{t}:=\mathbb{E}_{t}[Y\mid g(W)=1]$ Mean “coherence” rating assigned by domain experts to multi-turn conversational responses.
Conditional Variance $\theta_{t}:=\text{Var}_{t}[Y\mid g(W)=1]$ Variance in “helpfulness” ratings among novice programmers for code suggestions.
Conditional Quantile $\theta_{t}:=\inf\{y:P_{t}(Y\leq y\mid g(W)=1)\geq Q\}$ 90th percentile “safety” rating for high-risk queries flagged by content moderators.
@@
Estimation Goal. Given a sample of $N_{s}$ partial source observations $\mathcal{D}_{s}=\{(X^{s}_{j},V^{s}_{j},C^{s}_{j},C^{s}_{j}\cdot Y^{s}_{j},\hat{Y}^{s}_{j})\}_{j=1}^{N_{s}}$ and $N_{t}$ partial target observations $\mathcal{D}_{t}=\{(X^{t}_{i},V^{t}_{i},\hat{Y}^{t}_{i})\}_{i=1}^{N_{t}}$, our goal is to estimate a parameter summarizing system quality over $P_{t}$. [^2] We use superscripts $s$, $t$ on random variables to denote source and target membership. We omit these superscripts where the distribution is clear from context (e.g., $\mathbb{E}_{t}[Y]$ clearly refers to the target distribution). [^3] In the tuple $\mathcal{D}_{s}$, the shorthand $C^{s}_{j}\cdot Y^{s}_{j}$ denotes that ratings are only observed when $C^{s}_{j}=1$. As our running example in the main text and experiments, we consider the mean rating, $\theta_{t}:=\mathbb{E}_{t}[Y]$, which describes the average “safety” or “helpfulness” rating assigned by human raters to system outputs in the target distribution $P_{t}$.
@@ 3 Methodology
We now introduce our doubly-robust estimator for GenAI system quality estimation under evaluation sampling bias. The central challenge addressed by our approach is that our data is imperfect. While persona predictions are available, they may be a poor proxy for human ratings. Likewise, human ratings from the source distribution may suffer from evaluation sampling bias, leading to invalid estimates for the target parameter. We first introduce several naive approaches which might be used to tackle this problem ($\S$ 3.1). Then, we show that while each approach is insufficient in isolation, they can be combined to obtain valid coverage. ($\S$ 3.2). Our results presented in this section apply not only when $\theta_{t}:=\mathbb{E}_{t}[Y]$ (where they generalize Chernozhukov et al. (2023) to simultaneous covariate shift and selection bias) but also when $\theta_{t}$ is the solution to a generic M-estimation problem. Table 1 illustrates the range of statistical parameters our framework supports.
@@ 3.1 Baseline Approaches and Their Limitations
Re-weighting. Another approach is to re-weight samples from $P_{s}$ based on their probability of occurring under $P_{t}$. This approach requires correcting for covariate shift and selection bias in parallel, and does not use persona ratings. Formally, let $\omega_{0}(w)=\frac{dP_{t}}{dP_{s}}(w)$ denote the density ratio between $P_{t}(W)$ and $P_{s}(W)$, and let $\pi_{0}(w)=\mathbb{P}_{s}(C=1\mid W=w)$ denote the probability of rating completion. Under standard assumptions (see Appendix B), we have $\theta_{t}=\mathbb{E}_{s}[\alpha_{0}(W,C)Y]$, where $\alpha_{0}(W,C):=C\frac{\omega_{0}(W)}{\pi_{0}(W)}$. Thus, if one produces an ML estimate $\widehat{\alpha}$ of $\alpha_{0}$ (say by training models $\widehat{\omega}$, $\widehat{\pi}$ predicting $\omega_{0}$, $\pi_{0}$), they can compute an inverse propensity weighted (IPW) estimate $\hat{\theta}^{\text{ipw}}_{t}:=\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\hat{\alpha}(W^{s}_{j},C^{s}_{j})\cdot Y^{s}_{j}$. Again, estimates of $\alpha_{0}$ must converge at parametric rates in order to maintain coverage. Further, IPW suffers from high variance when propensities are small — a salient challenge when estimating system quality parameters over high-dimensional (e.g., text) data ($\S$ 4).
@@ 3.2 Doubly-Robust Estimator
where $\widehat{\mu}$ and $\widehat{\alpha}$ are estimates of $\mu_{0}$ and $\alpha_{0}$ that are assumed to be independent of the data. In Eq. 1, the left term evaluates the regression-based estimator over samples from the target distribution. Analogously to PPI++ (Angelopoulos et al., 2023b), this has the effect of using unlabeled data to reduce variance in the estimate. The right term corrects for bias in the human rating predictor $\hat{\mu}$ by re-weighting residualized source data to account for covariate shift and selection bias. This correction adjusts for bias in persona ratings via the residual term, and evaluation sampling bias via the learned re-weighting function $\hat{\alpha}$.
To construct confidence intervals, we also consider the variance estimate:
@@
where $\widehat{m}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\widehat{\mu}(W_{i}^{t},\hat{Y}_{i}^{t})$ and $\widehat{\gamma}$ is a scaling parameter (described in Algorithm 1). Since our mean and variance estimators require nuisance estimates that are independent of the data, we use $K$-fold cross-fitting to maximize efficiency (Chernozhukov et al., 2018). For each $k\leq K$, we train nuisance models on all data excluding the samples in fold $k$. We then use our nuisance estimates to produce a de-biased parameter estimate for the data in fold $k$. We finally average the per-fold parameter and variance estimates to maintain full data efficiency. See Algorithms 1 and 2 for full details.
@@
6:  Construct $\widehat{\theta}_{k}$ per Eq. 1 with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
7:  Construct $\widehat{\sigma}_{k}$ per Eq. 2 with $\widehat{\mu}:=\widehat{\mu}^{(-k)},\widehat{\alpha}:=\widehat{\alpha}^{(-k)}$, $\widehat{\gamma}:=\frac{N_{t}}{N_{s}}$, and samples $\mathcal{D}_{s,k}$ and $\mathcal{D}_{t}$.
@@
Assume the learner has access to samples $Z_{1}^{s},\dots,Z_{N_{s}}^{s}\sim P_{s}$ and $Z_{1}^{t},\dots,Z_{N_{t}}^{t}\sim P_{t}$ satisfying Assumptions 12 and the assumptions of Theorems B.2 and B.4 (all outlined in Appendix B). Then, letting $\widehat{\theta}$ and $\widehat{\sigma}^{2}$ denote the mean and variance returned by Algorithm 1, we have
@@
Theorem 3.1 relies on several key assumptions, formally outlined in Appendix B. Of these, the most important is double robustness, which requires the product of nuisance estimation errors to decay at parametric (i.e. $\sqrt{N_{t}}$) rates. Formally, this assumption can be expressed via the condition that
@@
and still maintain valid coverage. In other words, when we state that our estimator will provide valid coverage when either (i) a model trained to predict human ratings using persona ratings and source data observed under sampling bias ($\widehat{\mu}$), or (ii) a reweighting model that corrects for sampling bias ($\hat{\alpha}$) is of sufficient quality, we refer precisely to this product of errors condition (Eq. 3).
We also note that the above convergence rate does not directly depend on the quality of persona ratings. Rather, the persona ratings serve as an extra covariate onto which we can regress human ratings $Y$. When persona ratings are highly correlated with human ratings, we may obtain faster convergence rates for $\hat{\mu}$. However, this does not prohibit convergence even when the quality of persona ratings is low. This phenomenon is illustrated in our experiments below.
@@
To address this challenge, we leverage a “Riesz loss” (Chernozhukov et al., 2022a; b; 2023) to estimate $\alpha_{0}$. Rather than learning $\omega_{0}$ and $\pi_{0}$ independently, the Riesz loss directly learns the ratio $\alpha_{0}(w,c)$. For our setting, letting $\beta_{0}(w):=\omega_{0}(w)/\pi_{0}(w)$, the Riesz loss minimizer is given by:
@@
Therefore, to estimate $\alpha_{0}$, we minimize the finite-sample analogue of Eq. 4 using $\mathcal{D}_{s,k}^{c}$ and $\mathcal{D}_{t}$, then plug this into Algorithm 1 (see Appendix D for details). As we show in $\S$ 4, this Riesz loss approach significantly improves the quality of downstream estimates.